Could someone help me with this problem please:solving for x 7x=3
This an example of one I did: 3X=81
3X (3)^4
X=4 Thank you in advance Elle J

Answer by stanbon(75887)

please:solving for x 7x=3
This an example of one I did: 3X=81
3X (3)^4
X=4
--------------
Your example should be written as follows:
3^x = 81
3^x = 3^4
x = 4
--------------
Then your other problem is:
7^x = 3
xlog7 = log3
x = log3/log7
x = 1.771243....
---------------------
Cheers,
Stan H.